OpenStudy (anonymous):

The probability that a student passes mathematics class is 0.75, the probability that he passes history class is 0.80, and the probability that he passes mathematics and history is 0.60. Are the two events independent of each other?

OpenStudy (anonymous):

A. Yes, they are independent because P(M) ⋅ P(H) = P(M ∩ H) B. No, they are dependent because P(M) ⋅ P(H) = P(M ∩ H) C.Yes, they are independent because P(M) ⋅ P(H) ≠ P(M ∩ H) D. No, they are dependent because P(M) ⋅ P(H) ≠ P(M ∩ H)

OpenStudy (wilsondanielle):

You can multiply the probability of events if they are linked. Multiply the probabilities and see if it equals 0.6 to determine if they are dependent or independent. eg: flipping a coin gives a 1/2 chance of yielding heads. Flipping 2 coins gives a 1/4 (1/2x1/2) change of yielding two heads

OpenStudy (anonymous):

if \[P(M)\times P(H)=P(M\cap H)\] then they are independent if not , then they are dependent

OpenStudy (wilsondanielle):

Do you understand the notation used? (i.e. the "upside down u" and everything)

OpenStudy (anonymous):

I think so do i do .75 x .60=.45

OpenStudy (wilsondanielle):

no, use the probability of the individual subjects I.e. P(math) x P(history) = 0.75*0.8 = ?

OpenStudy (anonymous):

p(m)xP(H)=0.6

OpenStudy (wilsondanielle):

so p(m) x p(h) is equal to p(m+h). Does that make them dependent or independent?

OpenStudy (anonymous):

it makes them dependent?

OpenStudy (anonymous):

is \[.80\times .75=.60\]? it either is, or it is not

OpenStudy (anonymous):

it is .60

OpenStudy (anonymous):

so there are equal right?

OpenStudy (anonymous):

that makes the events "independent"

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so the answer would be Yes, they are independent because P(M) ⋅ P(H) = P(M ∩ H)

OpenStudy (wilsondanielle):

that's correct!

OpenStudy (anonymous):

ok thanks could u help me with one more?

OpenStudy (wilsondanielle):

sure!

OpenStudy (anonymous):

thx give me a sec

OpenStudy (anonymous):

There is a 70% chance that your car will get stuck in the snow during the next big snowfall. Given that you are already stuck in the snow with your car, the chance that you will require a tow truck to pull you out is 90%. What is the chance that you will get stuck in the snow with your car and will require a tow truck to pull you out?

OpenStudy (anonymous):

how would i solve this?

OpenStudy (wilsondanielle):

multiply the probabilities - IF you go out, there is a 0.7 chance you get stuck in the snow. IF you get stuck, there is a 0.9 chance you need a tow truck. The probability of both is like the coin example above, so just multiply them together.

OpenStudy (anonymous):

ok so it would be 0.63% ?

OpenStudy (wilsondanielle):

yep!

OpenStudy (anonymous):

could you check to see if i go this right i picked c One vase of flowers contains eight purple tulips and six yellow tulips. A second vase of flowers contains five purple tulips and nine yellow tulips. An example of dependent events is selecting a purple tulip from the first vase and then selecting a ____________ . . A. yellow tulip from the same vase, without replacing the first tulip B.purple tulip from the same vase, after replacing the first tulip C.yellow tulip from the second vase D. purple tulip from the second vase

OpenStudy (anonymous):

@wilsondanielle

OpenStudy (wilsondanielle):

which of these do you think depend on one another? a) same system, NOT reset b) same system, reset c) different system d) different system you're looking for something that is affected by the first action

OpenStudy (anonymous):

A?

OpenStudy (wilsondanielle):

A is the only one affected by the first tulip (because it is now missing, DECREASING the chance of drawing a purple tulip and INCREASING the chance of drawing a yellow tulip) so it would be correct

OpenStudy (anonymous):

so the answer would be a and not c?

OpenStudy (wilsondanielle):

yes. C is independent because it's in a different system - anything that happens in vase A doesn't affect Vase B because they aren't attached.

OpenStudy (anonymous):

ok thanks so much ur awesome

OpenStudy (wilsondanielle):

no problem!